• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Camsonne

    Log entry time 16:44:51 on July 15, 2006

    Entry number 174244

    This entry is a followup to: 174203

    keyword=Left VDC hot groups fixed

    Signals looked OK on scope. Hot groups were just misplugged ribbon cable shifted on the left, so only one pin for each channel was seeing signal. So no major problem for the VDC

    A copy of this log entry has been emailed to: bogdanw@jlab.org,segal@jlab.org